A sample contains both CO2 and Ne in unknown quantities.

Part A

If the sample contains a combined total of 1.45 mol and has a total mass of 51.0 g , what are the percentages of CO2 and Ne in the sample by mass?
Answer #1

Total Number of moles = 1-45 Total mars = 51 g. het mass of coz be eg Then mass of He = (51-x) g. : No. of moles of co = Mass